two number are in the ratio 3:8 and thei sum is 77.what are the number.

Answer by god2012(113)

Let 'x' be the common factor for the 2 numbers whose ratio is 3:8
Therefore, the numbers are 3x and 8x respectively.
Sum of the numbers = 77
or 3x + 8x = 77 or 11x = 77 or x = 7
3x = 3 * 7 = 21
8x = 8 * 7 = 56
the numbers are 21 and 56